Hernando county, Florida has an area of 589 square miles (there are 5,280 ft. in a mile). An inch of rain fell on the county. How many gallons of water fell on the county? A gallon of water weighs 81/3 pounds. A cubic foot of water weighs 62.4 pounds (its density). Write an equation for the number of gallons that fell in terms of the number of inches of rain. Let x be the number of inches of rain, let dw be the density of the water, let A be the area of the county, let wg be the weight of one gallon of water, and let N be the number of gallons of rain that fell on the county.
Answer by Boreal(15235)

Number of sq ft =5280 ft^2/mile*5280 ft^2/mile*589 sq miles=1.642 x 10^10
multiply by (1/12) for cubic feet since 1 inch of rain=1368364800 ft^3
62.4/(25/3)= number of gallons per cubic foot, so have 1.0246 x 10^10 gallons.
For rainwater harvesting, use 1000 ft^2 roof * 1 inch of rain, or 83.3333 ft^3 of water
62.4/(25/3) pounds/cubic foot/gallons=7.488 gallons per cubic foot, and that product, 624, is the number of gallons per 1000 ft^3 for an inch of rain.
x*(5280*5280*A)*62.4/[12*(25/3)]=N; x=inches of rain.